# PinRec: Outcome-Conditioned, Multi-Token Generative Retrieval for Industry-Scale Recommendation Systems

### v2

@@ 4.4. Serving
Figure 2. Serving flow for PinRec system (green boxes represent services, blue rectangles represent transmitted data, purple cylinders represent indexed data, and beige boxes are steps within the PinRec  Nvidia Triton ensemble).
PinRec is implemented with an in-house, optimized implementation of GPT-2, and served with NVidia Triton Inference Server’s ensemble setting and Python backend, to perform autoregressive generative retrieval. This multi-stage serving pipeline includes separate stages for fetching signals, featurizing Pins with learned ID embeddings, autoregressive inference for item generation, and Faiss retrieval (Figure 2).
@@ 4.4.1. Signal Fetching
Our Signal Service uses a Lambda architecture, which combines both batch and real-time processing. The batch signal is updated daily through a Spark pipeline that processes the past year’s positive user engagements and search queries, and stores the sequences and features for each user ID. The real-time component is backed by a RocksDB key-value store (Prasad, 2024) that allows us to efficiently fetch user engagements since the end time of the batch signal. The batch and real-time signals are de-duplicated and returned to PinRec. On the search surface, model inference also includes the search query issued by the user at request time as the final element in the sequence. To limit transport costs, all embeddings associated with sequence elements are quantized to INT8 precision. These are later de-quantized to FP16 for inference.
@@ 4.4.2. Learned ID Embeddings
As discussed in Section 4.1, our pin embedder includes an ID embedding table. Effectively, it is represented as a large in-memory table that maps each pin ID’s hashes to an embedding, followed by some MLP layers. As this ID embedder takes up significant space in memory, especially when replicated, it is separated out from the main PinRec model and runs in its own CPU memory-optimized service. It is served with torchscript to limit latencies to single-digit milliseconds.
4.4.3. Autoregressive Item Generation
To generate an ordered set of item recommendations, we employ autoregressive generation on the transformer with a KV Cache. Depending upon whether the model is outcome-conditioned or multi-token, we are able to sample or generate multiple representations respectively. After autoregressive generation, we use Faiss to retrieve a set of recommended items of size $N$ corresponding to the generated representations. Several optimizations are required for efficient inference:
@@ 4.4.3. Autoregressive Item Generation
Compression.
Given that we generate multiple output representations with autoregressive generation, we observe there are often embeddings that are similar to one another. During Faiss retrieval, these are likely to collapse into a similar set of output items. To optimize this, we “compress” each generated embedding into a prior uncompressed embedding, in the order it is generated, if it is more similar than a given cosine similarity threshold. Compression significantly helps reduce the median number of embeddings generated by each model, with an average ratio of 6:1 for an unconditioned PinRec model and a smaller ratio of 3:1 for the multi-token, outcome-conditioned variant.
Since we wish to fetch $N$ total items, we allocate these proportionally to each compressed embedding based on the total number of embeddings that were compressed into it. In the case of outcome-conditioning, we must perform this allocation independently for each action, in adherence with the chosen budget $\mathcal{B}$. The set of compressed embeddings alongside the allocations are passed to the Faiss retrieval stage.
@@ 4.4.4. Faiss Retrieval
Figure 3. Item recommendations from PinRec-OC for a user history with cat-related search queries and pin interactions. History in reverse chronological order. Additional examples in Section A.2.
Budget is: {repin: 0.4, outbound_click: 0.1, grid_click: 0.4, long_grid_click: 0.05, other: 0.05}. Generation runs for 6 steps (1 prefill, 5 decode).
Our nearest-neighbor retrieval involves a Faiss IVF-HNSW index running on CPU hosts, with inner product as the similarity metric. As with the main model, this is also served with NVidia Triton with batching enabled. We run $k$NN retrieval to fetch the allocated number of Pins for each embedding, with an overfetch factor for later de-duplication. Sample results are visualized for a manually specified user history in Figure 3.
@@ 5. Offline Evaluation
In this section, we dissect the performance of the proposed sequential recommendation system, PinRec,  on generative retrieval across a range of offline evaluations. Our experimental setting includes interactions on three major surfaces for the Pinterest platform, namely Homefeed, Related Pins, and Search. We demonstrate that outcome-conditioned generation allows for improved controllability and that multi-token generation leads to significantly improved diversity, performance, and efficiency.
@@ 5.1. Offline Evaluation Metrics
We construct the evaluation set using interactions for a set of users disjoint from the training set and evaluate whether the proposed technique can produce a representation to retrieve future item embeddings among a set of 1M random items (“index”). In such tasks, recall @ $k$ is a standard performance metric, defined as the proportion of target items correctly retrieved by the model within the top-$k$ ranked items across an index. However, standard recall assumes the existence of a single target per sample. As mentioned in Section 4.3.2, such an approach fails to capture the assumption of a lack of strict ordering between user activity within a small range of interactions $\Delta$. Further, when the goal of the retrieval model is candidate generation for a downstream ranking model, the task naturally has more than one unique plausible target, making standard recall less relevant.
We address these shortcomings by evaluating model performance on an unordered generalization of recall @ $k$. Given the autoregressively predicted item representations across all users $\mathbf{\hat{I}}$, target item embeddings across all users $\mathbf{I}$, and random item embeddings $\mathbf{N}$. Differently from standard recall, for each user $u\in\mathcal{U}$, unordered recall compares a set of predicted $\mathbf{\hat{I}}_{u}$ with a set of targets $\mathbf{I}_{u}$ of fixed size $\Delta$, calculating the proportion of targets $\mathbf{I}_{u}$ where there exists some $\mathbf{\hat{I}}_{u}$ that is within the top-$k$ ranked items. We formalize the notion of unordered recall in Equation 6.
(6) $\begin{split}\mathrm{ur}_{k}(\mathbf{\hat{I}},\mathbf{I},\mathbf{N})\coloneq\frac{1}{|\mathbf{\mathbf{\hat{I}}}|}\times\sum_{i=1}^{|\mathbf{\hat{I}}|}\mathbf{1}\left\{\min_{j,z\in[\Delta]}\left|\{\mathbf{n}\in\mathbf{N}\mid\mathbf{\hat{I}}_{ij}^{\top}\mathbf{n}\geq\mathbf{\hat{I}}_{ij}^{\top}\mathbf{I}_{iz})\}\right|\leq k\right\}\end{split}$
Note that since we perform autoregressive generation without teacher forcing, the transformer is not conditioned on any of the targets at any point.
Beyond unordered retrieval performance, we examine the diversity of the recommendations. After retrieving the top-$k$ predicted items for each user, we compute the proportion of unique items retrieved across the evaluation set. This allows us to quantify the degree to which learned representations are biased towards popular or frequently occurring Pins in the dataset.
@@ 5.2. Offline Results
In this section, we compare PinRec with an existing sequence-based user modeling representations, Pinnerformer (Pancha et al., 2022), SASRec (Kang and McAuley, 2018b), and TIGER (Rajput et al., 2023), in terms of offline evaluation metrics. Further, we quantify the impacts of our design on the performance and diversity of recommendations.
@@ 5.2.1. Comparison with Prior Method
To validate that PinRec outperforms existing sequential techniques in terms of item recommendation in offline evaluation, we compare its performance to PinnerFormer. We evaluate three different variants of PinRec.
@@
To ensure that the comparison between PinRec-{MT, OC} and other methods is fair, we adjust the number of autoregressive generation steps for multi-token generation to ensure that the total number of generated embeddings is held constant. As in training (where we cannot modify the evaluation dataset post-hoc), we condition on the target item’s action for outcome-conditioned variants.
Our comparison in Table 1 shows that PinRec outperforms a similarly sized Pinnerformer across the board. Importantly, outcome-conditioning demonstrates notable improvements over PinRec-UC (unconditioned) in unordered recall @ 10 (e.g., +2-3% on Homefeed and Related Pins). Above all, PinRec-{MT, OC} shows improvements over every method, with over +10% lift on Homefeed, over +20% on Related Pins, and +30% on search over PinRec-UC.
@@
TIGER (Rajput et al., 2023) 0.208 0.230 0.090
SASRec (Kang and McAuley, 2018b) 0.382 0.426 0.142
PinnerFormer (Pancha et al., 2022) 0.461 0.412 0.257
@@ 5.2.2. Outcome-Conditioned Generation
We seek to understand whether we can achieve “controllability” between the desired action and true action; for instance, does conditioning on a “repin” generate an output representation that truly aligns with content that will be repinned? If we are able to control the outcome through our item recommendations, then we will observe that for the specific action that we condition on, there are significant gains in offline recall. Conversely, we expect that the recall for items corresponding to other actions will decrease since our action-conditioned generation should ideally induce only that chosen action. As a baseline, we compare to the unconditioned variant of PinRec, which does not take into account the desired action.
For simplicity, we allocate the entire budget entirely to one action and perform offline evaluation conditioning across each action. We show the result of this analysis in Figure 4. Whenever we condition on an action outcome, we show notable gains over an unconditioned baseline for that action. Specifically, matching outcome conditioning yields 1.9% recall lift for true repins, 1.4% on true grid clicks, 4.8% on long grid clicks, and 6.2% on outbound clicks.
@@ 5.2.3. Utility of Multi-Token Generation
Beyond the improvement in performance, we evaluate the benefits of multi-token generation in terms of efficiency, diversity, and performance of generations relative to the number of chosen generations per timestep. We examine the tradeoff between generating more representations per timestep versus longer autoregressive generation, while keeping the number of output embeddings constant. Since the strongest offline performance aside from the multi-token variant is by PinRec-OC, we consider it as a baseline. For simplicity, we average the unordered recall @ 10 across the three surfaces.
We show the tradeoff between performance, diversity, and efficiency with the number of embeddings we generate in each generation step using multi-token generation in Figure 5. While next-token generation (i.e., 1 embedding per step) is the most inefficient and yields the poorest performance (where we perform similarly to PinRec-OC), we observe significant benefits across the board in increasing the number of embeddings per step. At 16 generations per step, we demonstrate $\sim$10x reduction in latency over PinRec-OC, with improved unordered recall (+16.0%) and diversity (+21.3%).
@@ 6. Online A/B Experiments
We perform online A/B experiments to evaluate PinRec as a candidate generator, measuring its impact on user engagement and confirming its online controllability. The items recommended by PinRec are passed through (unmodified) ranking and blending stages on each surface, alongside other candidate generators, to produce the final results shown to users. Since we employ a heterogeneous sequence of Pins and search queries, we focus on the Homefeed and Search surfaces. On these surfaces, we provide personalized item recommendations through PinRec, leading to significant lifts in action metrics, session metrics, and intent fulfillment.
@@ 6.1.2. Metrics
Pinterest uses “session fulfillment” as an informative and holistic whole-funnel metric to capture (a) the upper funnel of user intent expression (e.g., when a user clicks on a pin in the grid), (b) the middle funnel of curating Pins (e.g., repin), and (c) the lower funnel of realizing intent (e.g., clicking on a product pin to visit its website). Any of the aforementioned actions results in a session being marked as “fulfilled”, whereas the absence of these results in an “unfulfilled” session.
@@ 6.1.3. Results
In Figure 6, we visualize a Pareto front of the repin rate (relative to impressions) and grid click rate for the PinRec candidate generator by itself, with absolute numbers hidden due to non-public information. By reducing the grid click budget and correspondingly increasing the repin budget, we see the expected result as the grid click rate drops and repin rate increases (i.e., we demonstrate online “controllability”). Similarly, we observe a significant correlation between the budgets and the overall A/B experiment metric lifts for Homefeed grid clicks and repins in Figure 7. Further, we note that across both variants in Figure 7, for $\mathcal{B}(\text{repin})\geq 0.6$, there are statistically significant overall online lifts in Homefeed repin. We achieve statistically significant lifts in Homefeed grid click for all budget groups on each variant.
@@
(a) PinRec-OC
(b) PinRec-{MT, OC}
Figure 7. Visualization of Pareto fronts for online lift % in Homefeed grid clicks versus repin for PinRec variants.
To examine the holistic impact of our candidate generators, we examine the lift in fulfilled sessions for each PinRec variant, selecting the group which maximizes this metric. To demonstrate notable site-wide, session-level, and global metrics improvements, we present the CUPED-adjusted  (Deng et al., 2013) percentage lift in metrics across Homefeed and the entire site (Table 2) for the groups that maximize fulfilled sessions. Relative to the control group, the unconditioned variant of PinRec (PinRec-UC) demonstrates improvements in unfulfilled sessions and in Homefeed and site-wide grid clicks. In addition to those lifts, the outcome-conditioned variant (PinRec-OC) shows significant lift in site-wide fulfilled sessions. The multi-token variant, PinRec-{MT, OC}, significantly increased successful site sessions, time spent on site, and site-wide grid clicks, which we believe represent the most comprehensive site-wide improvements.
Table 2. CUPED-adjusted metrics improvements from online A/B experiments for PinRec variants as candidate generators in Homefeed.
@@ 6.2. Search
On the Search surface, we evaluate the unconditioned variant of PinRec. As shown in Table 3, PinRec-UC improves search metrics significantly relative to a control group. We observe session-wide search “fulfillment” increase, i.e., when a search session results in positive interaction such as repins or long outbound clicks.
Table 3. CUPED-adjusted metrics improvements from online A/B experiments for PinRec as candidate generator in Search retrieval.
@@
Overall Metrics Search Fulfillment Rate +0.48%
Action Metrics Search Repins +0.84%
Search Grid Clicks +1.46%
@@ 7. Conclusion
We introduced PinRec, a generative retrieval system developed for Pinterest’s Homefeed and Search feed. To the best of our knowledge, this paper represents the first comprehensive study on deploying generative retrieval in a web-scale application. The incorporation of two key innovations — outcome-conditioned generation and multi-token inference — enabled PinRec to enhance several business metrics without substantial cost increases.
Several promising future directions can be pursued. First, applying generative approaches to ranking, as explored by Zhai et al. (Zhai et al., 2024), could be beneficial. Second, optimizing input sequences to the model, akin to prompt engineering in large language models, is another avenue worth exploring. Third, developing language model alignment techniques similar to Ouyang et al. (Ouyang et al., 2022) could enhance the use of engagement signals from users. Overall, this paper highlights a promising avenue for utilizing generative retrieval in large-scale recommender systems.
Acknowledgements.
We thank collaborators such as Deepak Agarwal, Divyansh Agarwal, Edoardo Botta, Bee-Chung Chen, Bowen Deng, Pong Eksombatchai, Kurchi Subhra Hazra, Se Won Jang, Saurabh Vishwas Joshi, Krishna Kamath, Sujay Khandagale, James Li, Cristian Lopez, Matt Madrigal, Lei Pan, and Ruijia Wang for their critical contributions to our work and for facilitating productive discussions.

### v4

@@ 4.5. Serving
Figure 3. Serving flow for the PinRec system. Green boxes represent services, blue rectangles represent transmitted data, purple cylinders represent indexed data, and beige boxes are steps within the PinRec NVIDIA Triton ensemble.
PinRec is implemented with an in-house, optimized implementation of transformer and served with NVIDIA Triton Inference Server’s ensemble setting and Python backend to perform autoregressive generative retrieval. This multi-stage serving pipeline includes separate stages for fetching signals, featurizing Pins with learned ID embeddings, autoregressive inference for item generation, and Faiss retrieval (Figure 3).
@@ 4.5.1. Signal Fetching
Our Signal Service uses a Lambda architecture that combines both batch and real-time processing. The batch signal is updated daily through a Spark pipeline that processes the past year’s positive user engagements and search queries, storing the sequences and features for each user ID. The real-time component is backed by a RocksDB key-value store (kvstore) that efficiently retrieves user engagements occurring after the batch signal’s cutoff time. The batch and real-time signals are deduplicated and returned to PinRec. On surfaces with context (e.g., Search, Related Pins), model inference also includes the context (e.g., search query, closeup Pin) issued by the user at request time as the final element in the sequence. To limit transport costs, all embeddings associated with sequence elements are quantized to INT8 precision and later dequantized to FP16 for inference.
@@ 4.5.2. Learned ID Embeddings
As discussed in Section 4.2, our pin embedder includes an ID embedding table that maps each pin ID’s hash to an embedding, followed by MLP layers. Since this ID embedder requires significant memory, especially when replicated, it runs as a separate CPU memory-optimized service to avoid competing with the main model for GPU resources. It is served with torchscript to achieve single-digit millisecond latencies.
4.5.3. Autoregressive Inference
We serve the main PinRec model for autoregressive generation on NVIDIA L40S GPUs. To generate an ordered set of item recommendations efficiently, we employ several optimizations for latency and cost efficiency:
@@ 4.5.4. Faiss Retrieval
Figure 4. Item recommendations from PinRec-OC for a user history with cat-related search queries and pin interactions (history shown in reverse chronological order). Additional examples are provided in Section A.2.
Our nearest-neighbor retrieval involves a Faiss IVF-HNSW index running on CPU hosts, with inner product as the similarity metric. As with the main model, this is also served with NVIDIA Triton with batching enabled. We run kkNN retrieval to fetch the allocated number of Pins for each embedding, with an overfetch factor for later de-duplication.
Sample results from this complete serving pipeline are visualized for a manually specified user history in Figure 4.
@@ 5. Offline Evaluation
In this section, we dissect the performance of the proposed generative recommendation system, PinRec, on generative retrieval across a range of offline evaluations. For the results presented, PinRec uses a 12-layer transformer model (12 attention heads, 768 hidden dimensions) with an 80M-row ID embedding table (256 total dimensions, 8 dimensions per hash). Our experimental setting includes interactions on three major surfaces for the Pinterest platform, namely Homefeed, Related Pins, and Search. We demonstrate that outcome-conditioned generation allows for improved controllability and that multi-token generation leads to significantly improved diversity, performance, and efficiency.
@@ 5.1. Offline Evaluation Metrics
We construct the evaluation set using interactions from users disjoint from the training set and evaluate whether the proposed technique can produce representations that retrieve future item embeddings from a set of $1M$ random items (the ”index”). For such retrieval tasks, recall@$k$ is a standard performance metric, measuring the proportion of target items correctly retrieved within the top-$k$ ranked items.
We evaluate model performance using a sequence generalization of recall@$k$ that accounts for the multi-embedding nature of our autoregressive generation. Unlike standard recall, which compares individual predictions to individual targets, our unordered recall metric compares sets of predictions to sets of targets for each user.
Given output item representations $\hat{\mathbf{I}}_{u}=\{\hat{\mathbf{i}}_{u,1},\hat{\mathbf{i}}_{u,2},\ldots,\hat{\mathbf{i}}_{u,m}\}$ and target item embeddings $\mathbf{I}_{u}=\{\mathbf{i}_{u,1},\mathbf{i}_{u,2},\ldots,\mathbf{i}_{u,n}\}$ for user $u$, along with random negative embeddings $\mathbf{N}$, we calculate the proportion of target items for which at least one predicted embedding retrieves it within the top-$k$ results:
$\begin{split}\mathrm{ur}_{k}(\mathbf{\hat{I}},\mathbf{I},\mathbf{N})\coloneq\frac{1}{|\mathbf{\mathbf{\hat{I}}}|}\times\sum_{i=1}^{|\mathbf{\hat{I}}|}\mathbf{1}\left\{\min_{j,z\in[\Delta]}\left|\{\mathbf{n}\in\mathbf{N}\mid\mathbf{\hat{I}}_{ij}^{\top}\mathbf{n}\geq\mathbf{\hat{I}}_{ij}^{\top}\mathbf{I}_{iz})\}\right|\leq k\right\}\end{split}$
where $\text{rank}(\mathbf{i},\hat{\mathbf{i}},\mathbf{N})$ denotes the rank of target item $\mathbf{i}$ when items in $\mathbf{N}\cup{\mathbf{i}}$ are sorted by cosine similarity to prediction $\hat{\mathbf{i}}$. This metric captures whether any of the generated embeddings successfully retrieves each target item, reflecting the practical scenario where multiple predictions can contribute to successful recommendation.
We perform autoregressive generation without teacher forcing, ensuring the transformer is not conditioned on any target items during prediction, making the evaluation more realistic for deployment scenarios.
Beyond retrieval performance, we examine recommendation diversity by computing the proportion of unique items retrieved across all users in the evaluation set after selecting top-$k$ predictions. This quantifies whether learned representations exhibit bias toward popular items versus providing diverse recommendations across the user base.
@@ 5.2. Offline Results
In this section, we compare PinRec against existing sequence-based user modeling approaches, including Pinnerformer (Pinnerformer22), SASRec (kang2018self), and the leading generative recommendation method TIGER (Tiger23). We also evaluate our transformer architecture against the proposed HTSU (HSTU24) module. Furthermore, we quantify the impacts of our design choices on both recommendation performance and diversity.
@@ 5.2.1. Comparison with Prior Method
We evaluate three different variants of PinRec.
@@
To ensure that the comparison between PinRec-{MT, OC} and other methods is fair, we adjust the number of autoregressive generation steps for multi-token generation to ensure that the total number of generated embeddings is held constant. As in training (where we cannot modify the evaluation dataset post-hoc), we condition on the target item’s action for outcome-conditioned variants. For simplicity, we reproduce results for the outcome-conditioned variant only when comparing against HTSU (additional details in Section A.1.1). The impact of negative sampling and model scaling effects are analyzed in Section A.1.2 and Section A.1.3, respectively.
Our comparison in Table 1 shows that PinRec outperforms other baselines across the board. Importantly, outcome-conditioning demonstrates notable improvements over PinRec-UC (unconditioned) in unordered recall @ 10 (e.g., +2-3% on Homefeed and Related Pins). Above all, PinRec-{MT, OC} shows improvements over every method, with over +10% lift on Homefeed, over +20% on Related Pins, and +30% on search over PinRec-UC.
@@
SASRec (kang2018self) 0.382 0.426 0.142
PinnerFormer (Pinnerformer22) 0.461 0.412 0.257
TIGER (Tiger23) 0.208 0.230 0.090
HTSU-OC (HSTU24) 0.596 0.539 0.179
@@ 5.2.2. Outcome-Conditioned Generation
We assess whether conditioning on a desired action (e.g., “repin”) leads to outputs that elicit that action. If our recommendations successfully control outcomes, we expect higher offline recall for the conditioned action and lower recall for others. As a baseline, we compare against the unconditioned PinRec variant.
For simplicity, we allocate the entire budget to one action and perform offline evaluation conditioning across each action. We show the result of this analysis in Figure 5. Whenever we condition on an action outcome, we show notable gains over an unconditioned baseline for that action. Specifically, matching outcome conditioning yields 1.9% recall lift for true repins, 1.4% on true grid clicks, 4.8% on long grid clicks, and 6.2% on outbound clicks.
@@ 5.2.3. Utility of Multi-Token Generation
We evaluate the benefits of multi-token generation as a function of the number of generations produced per timestep. Specifically, we analyze the tradeoff between generating more representations in fewer steps versus fewer representations over more steps, while keeping the total number of output embeddings constant. Since the strongest offline performance aside from the multi-token variant is by PinRec-OC, we consider it as a baseline. For simplicity, we average the unordered recall @ 10 across the three surfaces.
We show the tradeoff between performance, diversity, and efficiency with the number of embeddings we generate in each generation step using multi-token generation in Figure 6. While next-token generation (i.e., 1 embedding per step) is the most inefficient and yields the poorest performance (where we perform similarly to PinRec-OC), we observe significant benefits across the board in increasing the number of embeddings per step. At 16 generations per step, we demonstrate a $\sim$10x reduction in latency over PinRec-OC, with improved unordered recall (+16.0%) and diversity (+21.3%).
@@ 6. Online A/B Experiments
We perform online A/B experiments to evaluate PinRec as a candidate generator, measuring its impact on user engagement and confirming its online controllability. The items recommended by PinRec are passed through (unmodified) ranking and blending stages on each surface, alongside other candidate generators, to produce the final results shown to users. We tested on the three major surfaces: Homefeed, Search and Related Pins.
@@ 6.1.2. Metrics
We use engagement actions such as repins, grid clicks, and time spent. Additionally, we count “fulfilled sessions,” where the user takes at least one action in the session, and “unfulfilled sessions,” where the user does not take any action.
@@ 6.1.3. Results
In Figure 7, we visualize a Pareto front of the repin rate (relative to impressions) and grid click rate for the PinRec candidate generator by itself, with absolute numbers hidden due to non-public information. By reducing the grid click budget and correspondingly increasing the repin budget, we see the expected result as the grid click rate drops and repin rate increases. Similarly, we observe a significant correlation between the budgets and the overall A/B experiment metric lifts for Homefeed grid clicks and repins in Figure 8 in Appendix.
@@
For each PinRec variant, we selected the variant with the best fulfilled sessions and measured the percentage lift in metrics across Homefeed and the entire site (Table 2). Relative to the control group, the unconditioned variant of PinRec, PinRec-UC, demonstrates improvements in unfulfilled sessions and in Homefeed and site-wide grid clicks. In addition to those lifts, the outcome-conditioned variant, PinRec-OC, shows significant lift in site-wide fulfilled sessions. The multi-token variant, PinRec-{MT, OC}, significantly increased successful site sessions, time spent on site, and site-wide grid clicks.
Table 2. A/B experiments for PinRec variants as candidate generators in Homefeed.
@@ 6.2. Search
On the Search surface, we evaluate the unconditioned variant of PinRec. As shown in Table 3, PinRec-UC improves search metrics significantly relative to a control group, with increased session-wide search fulfillment (sessions resulting in positive interactions like repins or long clicks). When deployed for promoted product retrieval, the same embeddings achieve a 1.83% reduction in CPA and 1.87% increase in shopping conversion volume in the US, indicating improved relevance and user engagement with sponsored content.
Table 3. A/B experiments for PinRec as candidate generator in Search retrieval.
@@
Overall Metrics Search Fulfillment Rate +2.27%
Action Metrics Search Repins +4.21%
Search Grid Clicks +3.00%
6.3. Related Pins
On the Related Pins surface, we evaluate the outcome-conditioned variant of PinRec. As shown in Table 4, PinRec-OC improves Related Pins metrics significantly relative to a control group. We observe session-wide increase in fulfilled sessions and increase in sitewide engagement.
Table 4. A/B experiments for PinRec variants as candidate generators in Related Pins.
Type Metric PinRec-OC
Overall Fulfilled Sessions +0.26%
Metrics Time Spent +0.30%
Unfulfilled Sessions -1.07%
Action Site-wide Grid Clicks +0.50%
Metrics Sitewide Repins +0.63%
@@ 7. Conclusion
We introduced PinRec, a generative retrieval system developed for Pinterest’s Homefeed, Search and Related Pins feed. To the best of our knowledge, this paper represents the first comprehensive study on deploying generative retrieval in a web-scale application. The incorporation of two key innovations — outcome-conditioned generation and temporal multi-token inference — enabled PinRec to enhance several business metrics without substantial cost increases.
Several promising future directions can be pursued. First, applying generative approaches to ranking, as explored by Zhai et al. (HSTU24), could be beneficial. Second, optimizing input sequences to the model, akin to prompt engineering in large language models, is another avenue worth exploring. Third, developing language model alignment techniques similar to Ouyang et al. (PPO22) could enhance the use of engagement signals from users. Overall, this paper highlights a promising avenue for utilizing generative retrieval in large-scale recommender systems.